# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 6 Numerical Experiments
We evaluate the performance of the norm test and (augmented) inner product test with AdaGrad(-Norm) and SGD in image classification tasks, employing logistic regression and CNNs on the MNIST (LeCun et al., 1998) and CIFAR-10 (Krizhevsky, 2009) datasets. We remark that training of larger models often requires multiple workers and data parallelism, such as Distributed Data Parallel (DDP) (Li et al., 2020) and Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023). However, adapting our proposed adaptive batch size schemes for data parallelism presents additional complexities and remains an area for future research. For this study, we concentrate on smaller models and datasets, utilizing a single GPU without parallelism, with the goal of demonstrating the concept rather than achieving state-of-the-art results. Following the observation in Bollapragada et al. (2018a), we exclude the orthogonality test from our implementation to minimize computational burden. Given the typically large number of parameters $d$, conducting coordinate-wise norm tests for AdAdaGrad is not computationally practical, hence the standard norm test is applied.
For the implementation of the proposed schemes, per-sample gradients are computed using JAX-like composable function transforms called torch.func in PyTorch 2.0+. Numerical experiments are carried out on workstations with NVIDIA RTX 2080Ti 11GB (for MNIST) or A100 40/80GB (for CIFAR-10) GPUs, with PyTorch 2.1.0 (Paszke et al., 2019) and Lightning Fabric 2.1.3 (Falcon and The PyTorch Lightning team, 2019). The AdaGrad-Norm implementation is taken from Ward et al. (2019). Details of the experiments, such as the hyperparameter setting, can be found in Appendix C.
6.1 Multi-class Logistic Regression on MNIST
We initially applied our methods to a ten-class logistic regression problem using the MNIST dataset, a scenario representing a smooth convex challenge. Our experiments were conducted with an equal training budget of 6 million samples (equivalent to 100 epochs), setting a maximum batch size of 60,000 (i.e., the full batch) for all approaches. To highlight the adaptivity and flexibility of our proposed methods, we refrained from conducting an exhaustive search for optimal values of $\alpha$, $\eta$, and $\vartheta$, and we did not employ any learning rate schedules across the methods. The outcomes, including the number of iterations required (steps), average batch sizes (bsz.), final training loss (loss), and final validation accuracy (acc.), are documented in Table 1.
Table 1: Multi-class logistic regression on MNIST
@@ 6.1 Multi-class Logistic Regression on MNIST
AdaSGD $\eta=0.10$ 351 17131 1.04 0.82
AdaSGD $\eta=0.25$ 1029 5831 0.67 0.86
AdAdaGrad-Norm $\eta=0.10$ 596 10060 1.50 0.78
AdAdaGrad-Norm $\eta=0.25$ 2462 2437 1.02 0.83
AdAdaGrad $\eta=0.10$ 126 47282 0.54 0.88
AdAdaGrad $\eta=0.25$ 274 21918 0.46 0.90
AdaSGD $\vartheta=0.05$ 717 8362 0.76 0.85
AdaSGD $\vartheta=0.10$ 1804 3327 0.54 0.88
AdAdaGrad-Norm $\vartheta=0.05$ 1127 5322 1.28 0.80
AdAdaGrad-Norm $\vartheta=0.10$ 5349 1122 0.80 0.85
The results, coupled with the graphical representations in Figure 1, reveal that AdAdaGrad, utilizing the norm test with $\eta=0.25$, outperforms others in terms of both training loss and validation accuracy. These findings underscore the importance of adaptive (and coordinate-wise) learning rates within adaptive sampling methods. Despite the best performance of AdAdaGrad using the norm test with $\eta=0.25$, we note that the one with $\eta=0.10$ is more hardware-efficient—having an average batch size of more than 47,000, totaling only 126 steps (i.e., gradient evaluation) to run through all 6 million samples. This implies an intrinsic trade-off between computationally efficiency and generalization. Interestingly, AdAdaGrad-Norm, while theoretically simpler to analyze, might underperform compared to AdaSGD for this specific convex problem and set of hyperparameters. It is also noteworthy that the norm test tends to increase batch sizes more aggressively than the inner product test, leading to more efficient utilization of available GPU memory.
6.2 Three-layer CNN on MNIST
@@
For the same image classification problem, we also experiment with a three-layer CNN with a nonconvex objective. We use the same training budget and maximum batch size. In Figure 2 (see Appendix C), similar to the previous convex problem, AdAdaGrad also outperforms AdAdaGrad-Norm and AdaSGD by a certain margin. However, unlike the logistic regression problem, AdAdaGrad-Norm enjoys similar performance to AdaSGD in terms of validation accuracy, in spite of slightly higher training loss. We also observe from Table 4 that AdAdaGrad using the norm test with $\eta=0.1$ is able to achieve a validation accuracy of 96% with only 148 iterations with an average batch size of more than 40,000. It makes use of full batch (i.e., full gradient without batch sampling) for the last 70% of its training budget, taking full advantage of available GPU memory.
6.3 Three-layer CNN on CIFAR-10
7 Conclusion and Future Work
@@
In this work, we showcase the versatility of adaptive sampling methods as generic adaptive batch size schemes suitable for adaptive gradient optimizers, supported by convergence guarantees and numerical results. This revelation opens up a myriad of intriguing research directions for adaptive batch size schemes as a promising and flexible option for large-scale model training. On the theoretical front, delving deeper into the convergence guarantees of this class of methods when combined with other stochastic gradient optimizers, such as momentum-based methods, as well as proximal or projection methods for constrained problems with deterministic nonsmooth regularizers, would be fascinating. Practically, exploring the implementation of adaptive batch size schemes under various paradigms of parallelism for large-scale distributed training—including data, tensor, and pipeline parallelism (Shoeybi et al., 2019; Rajbhandari et al., 2020; Smith et al., 2022; Zhao et al., 2023)—is worthwhile. This exploration aims to ensure these schemes are viable for large-scale applications, such as the (pre-)training of autoregressive language and image models. Additionally, examining the impact of adaptive batch size schemes for adaptive gradient methods, in contrast to those designed for SGD, particularly for language models based on transformers in addition to the CNNs for vision tasks discussed in this paper, is crucial. This is because, unlike the marginal utility of adaptive methods for CNNs and RNNs (Wilson et al., 2017), adaptive gradient methods like Adam significantly outperform SGD in optimizing transformers (Zhang et al., 2020b, c; Jiang et al., 2022; Kunstner et al., 2023; Pan and Li, 2023; Ahn et al., 2023).
@@ Acknowledgments
The research of Han Liu is supported by NIH R01LM01372201, NSF RI 1840857, NSF TRIPOD 1740735, NSF DMS1454377-CAREER, NSF IIS 1546482, along with an Alfred P. Sloan Fellowship. The research of Mladen Kolar is supported in part by NSF ECCS-2216912.
@@
Appendix A The Overall Algorithm
@@
$\hskip 22.76219pt\mathsf{T}_{j}(x_{k};\mathcal{B}_{k},\eta)\coloneqq\frac{1}{b_{k}-1}\sum_{i\in\mathcal{B}_{k}}\left(\partial_{j}f(x_{k};\xi_{i})-\partial_{j}F_{\mathcal{B}_{k}}(x_{k})\right)^{2}\big{/}\left(\eta^{2}(\partial_{j}F_{\mathcal{B}_{k}}(x_{k}))^{2}\right),\;j\in\llbracket d\rrbracket$
@@
                      $\mathsf{T}_{\mathrm{ip}}(x_{k};\mathcal{B}_{k},\vartheta)\coloneqq\mathrm{Var}_{i\in\mathcal{B}_{k}}(\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle)\big{/}\left(\vartheta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{4}\right)$
@@
Compute the orthogonality test statistic $\mathsf{T}_{\mathrm{ortho}}(x_{k};\mathcal{B}_{k},\nu)\coloneqq\mathsf{V}(x_{k};\mathcal{B}_{k})/\left(\nu^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}\right)$
@@
Appendix B Proofs of Main Text
@@ B.1 Proof of Equation 3
Note that $-\nabla F_{\mathcal{B}}(x)$ is a descent direction for $F$ if and only if $\langle\nabla F_{\mathcal{B}}(x),\nabla F(x)\rangle\geqslant 0$. Then we have, for $\eta\in\left[0,1\right[$,
$(\forall x\in\mathbb{R}^{d})\quad\|\nabla F_{\mathcal{B}}(x)-\nabla F(x)\|^{2}=\|\nabla F_{\mathcal{B}}(x)\|^{2}-2\langle\nabla F_{\mathcal{B}}(x),\nabla F(x)\rangle+\|\nabla F(x)\|^{2}\leqslant\eta^{2}\|\nabla F(x)\|^{2}\\ \iff-2\langle\nabla F_{\mathcal{B}}(x),\nabla F(x)\rangle\leqslant-(1-\eta^{2})\|\nabla F(x)\|^{2}-\|\nabla F_{\mathcal{B}}(x)\|^{2}\leqslant 0,$
@@
B.2 Preparatory Definitions, Propositions and Lemmas
@@
B.2.1 Formal Statements of Proposition 5.1
@@
Proposition B.1 (Exact variance norm test).
Suppose that the condition of the exact variance norm test (6) is satisfied for every iteration $k\in\mathbb{N}^{*}$, i.e., for some constant $\eta\in\mathbb{R}_{++}$,
@@
$\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant(1+\eta^{2})\|\nabla F(x_{k})\|^{2}.$ (B.1)
@@
Proposition B.2 (Exact variance inner product test and orthogonality test).
Suppose that the conditions of the exact variance inner product test (7) and the exact variance orthogonality test (9) are satisfied for every iteration $k\in\mathbb{N}^{*}$, i.e., for some constant , i.e., for some constant ϑ++italic-$\vartheta\in\mathbb{R}_{++}$,
@@
$\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant(1+\vartheta^{2}+\nu^{2})\|\nabla F(x_{k})\|^{2}.$ (B.2)
@@
This is part of the result of Bollapragada et al. (2018a, Lemma 3.1). We include its proof here with our notation for completeness.
@@
$\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant\frac{\mathbb{E}_{k}\left[\left(\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle\right)^{2}\right]}{\|\nabla F(x_{k})\|^{2}}+\nu^{2}\|\nabla F(x_{k})\|^{2}.$ (B.3)
It remains to bound the first term on the right hand side of (B.3), using the condition of the exact variance inner product test. Since (7) is satisfied, the following inequality holds:
@@
Along with (B.3), this yields
@@
which is identical to (6). The desired conclusion follows from Proposition B.1. ∎
B.2.3 Technical Lemmas

### v2

@@ 6 Numerical Experiments
We evaluate the performance of the norm test and the (augmented) inner product test with AdaGrad(-Norm) and SGD for image classification, employing logistic regression (Section D.1) and a three-layer CNN on the MNIST dataset (LeCun et al., 1998), and a three-layer CNN (Section D.3) and ResNet-18 (He et al., 2016) on the CIFAR-10 dataset (Krizhevsky, 2009). We note that training larger models often requires multiple workers and data parallelism, such as Distributed Data Parallel (DDP) (Li et al., 2020) and Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023). Extending and implementing our proposed schemes under data parallelism presents additional complexities and remains an area for future research. We thus concentrate on smaller models and datasets with the goal of demonstrating the concept rather than achieving state-of-the-art results. Given the typically large number of parameters $d$, conducting coordinate-wise norm tests for AdAdaGrad is not computationally practical, so the standard norm test is applied. Details of the experiments, such as the hyperparameter setting, and additional results can be found in Appendix D.
6.1 Three-layer CNN on MNIST
Table 1: Three-layer CNN on MNIST
@@ 6.1 Three-layer CNN on MNIST
SGD N/A 366 16384 0.51 0.87
SGD N/A 183 32768 1.54 0.75
AdaGrad N/A 199 30000 0.10 0.97
AdaGrad N/A 149 40000 0.13 0.96
AdaSGD $\eta=0.10$ 256 23546 0.79 0.83
AdaSGD $\eta=0.25$ 383 15627 0.48 0.88
AdAdaGrad-Norm $\eta=0.10$ 226 26567 0.88 0.83
AdAdaGrad-Norm $\eta=0.25$ 435 13830 0.54 0.87
AdAdaGrad $\eta=0.10$ 149 40057 0.15 0.96
AdAdaGrad $\eta=0.25$ 198 30152 0.13 0.97
@@
We first train a three-layer CNN on the MNIST classification problem with a nonconvex objective. Our experiments were carried out with an equal training budget of 6 million samples (equivalent to 100 epochs), setting a maximum batch size of 60,000 (i.e., the full batch) for all approaches. We can measure training efficiency based on the number of (gradient) steps, instead of wall-clock time (available in Table 5), which is device dependent (Shallue et al., 2019). Adaptive batch size methods begin from a small batch size 8 and gradually increase to the maximum possible batch size of 60000 (full batch). In Figure 1, AdAdaGrad outperforms AdAdaGrad-Norm and AdaSGD in terms of validation accuracy (generalization) by a certain margin. AdAdaGrad-Norm enjoys similar performance to AdaSGD, in spite of slightly higher training loss. We also observe from Table 1 that AdAdaGrad using the norm test with $\eta=0.1$ is able to achieve a validation accuracy of 96% with only 149 iterations with an average batch size of more than 40,000. It uses full batches for the last 70% of its training budget, taking full advantage of the available GPU memory. Referring to Table 5 in Section D.2, we observe the generalization gap between small constant batch sizes and large batch sizes, while using small batch sizes requires substantially longer training time and a higher number of steps, leading to lower training efficiency. Our proposed methods are capable of balancing this inevitable trade-off between training efficiency and generalization by introducing adaptive batch size schemes, without the need of extensive tuning of learning rates or pre-specified learning rate schedules.
Figure 1: Training loss, validation accuracy and batch sizes of AdaSGD, AdAdaGrad and AdAdaGrad-Norm for a three-layer CNN on the MNIST dataset.
6.2 ResNet-18 on CIFAR-10
We then train a ResNet-18 for image classification on the harder CIFAR-10 dataset. We use a training budget of 10 million samples (200 epochs) and a maximum batch size of 50,000 samples. Although the focus of this work is on AdaGrad, we also empirically study the effect of adaptive batch size schemes for Adam attributed to its ubiquity in deep learning, which is known as AdAdam.
AdAdaGrad.
In Figure 2, we observe that we need to choose a rather small $\eta$ to make use of full batches for training at later stages, when this larger model is trained. Comparing AdaGrad with a constant batch size of 50000 and AdAdaGrad with $\eta=0.025$ (see also Table 2), AdAdaGrad is able to use full batches in most of the later training stages, while achieving high accuracy with 23 more steps. In general (see also Table 10 in Section D.4),our proposed scheme is able to close the generalization gap between smaller and larger constant batches (e.g., the curves of $\eta=0.075$ and $0.1$ lying in the gap between those of constant batch sizes 4096 and 8192).
Figure 2: AdaGrad and AdAdaGrad for ResNet-18 on the CIFAR-10 dataset.
Figure 3: Adam and AdAdam for ResNet-18 on the CIFAR-10 dataset.
AdAdam.
In Figure 3, we observe a similar trend of AdAdam to AdAdaGrad but with faster convergence and larger batch size increases. It is worth noticing from Table 2 that AdAdam with $\eta=0.1$ and an average batch size of 8880 outperforms Adam with a smaller constant batch size of 8192 in terms of validation accuracy, while requiring almost 100 fewer steps. This indicates that our proposed scheme is even more beneficial when used with Adam.
Table 2: ResNet-18 on CIFAR-10
Scheme test steps bsz. loss acc.
AdaGrad N/A 305 32768 0.9684 0.5816
AdaGrad N/A 199 50000 1.3625 0.4708
Adam N/A 1220 8192 0.0004 0.8946
Adam N/A 610 16384 0.0028 0.8628
AdAdaGrad $\eta=0.025$ 222 44934 1.2770 0.5107
AdAdaGrad $\eta=0.05$ 485 20615 0.6204 0.7079
AdAdam $\eta=0.05$ 426 23463 0.0061 0.8228
AdAdam $\eta=0.1$ 1126 8880 0.0000 0.9042
7 Concluding Remarks
@@
In this work, we demonstrate the versatility of adaptive sampling methods as generic adaptive batch size schemes suitable for adaptive gradient optimizers, supported by convergence guarantees and numerical results. This revelation opens up a myriad of intriguing research directions for adaptive batch size schemes as a promising and flexible option for large-scale model training. On the theoretical front, delving deeper into the convergence guarantees of this class of methods when combined with other stochastic gradient optimizers, such as momentum-based methods, as well as proximal SGD methods for constrained problems with deterministic nonsmooth regularizers, would be fascinating. Practically, exploring the implementation of adaptive batch size schemes under various parallelism paradigms for large-scale distributed training—including data, tensor, and pipeline parallelism (Shoeybi et al., 2019; Rajbhandari et al., 2020; Smith et al., 2022; Zhao et al., 2023)—is worthwhile. This exploration aims to ensure these schemes are viable for large-scale applications, such as the (pre-)training of autoregressive language and image models. Furthermore, examining the impact of adaptive batch size schemes for adaptive gradient methods, in contrast to those designed for SGD, particularly for language models based on transformers in addition to CNNs for vision tasks discussed in this paper, is crucial. This is because, unlike the marginal utility of adaptive methods for CNNs and RNNs (Wilson et al., 2017), adaptive gradient methods such as Adam significantly outperform SGD in optimizing transformers (Zhang et al., 2020b, c; Jiang et al., 2022; Kunstner et al., 2023; Pan and Li, 2023; Ahn et al., 2023).
@@ Acknowledgments
The research of Han Liu is supported by NIH R01LM01372201, NSF RI 1840857, NSF TRIPOD 1740735, NSF DMS1454377-CAREER, NSF IIS 1546482, along with an Alfred P. Sloan Fellowship. The research of Mladen Kolar is supported in part by NSF ECCS-2216912. This research is supported in part through the computational resources and staff contributions provided for the Data Science Institute at the University of Chicago, through its AI + Science Research Funding Initiatives.
@@
Appendix A Further Related Work
A.1 Adaptive Sampling Methods
Adaptive sampling methods have also been extended to tackle problems outside the unconstrained optimization setting. Xie et al. (2023) proposed a proximal extension of the norm test and the inner product test to minimize a convex composite objective, which comprises a stochastic function and a deterministic, potentially nonsmooth function. Beiser et al. (2023) explored deterministic constrained problems, including those with nonconvex objectives. Moreover, adaptive sampling methods have been applied to various optimization problems and algorithms, such as (Bollapragada et al., 2018b) for L-BFGS, (Berahas et al., 2022) for sequential quadratic programming (SQP) in the context of equality-constrained problems, (Bollapragada et al., 2023) for augmented Lagrangian methods, and (Bollapragada and Wild, 2023) for quasi-Newton methods.
A.2 Convergence of Stochastic Gradient Methods
We give a brief overview on the convergence results of SGD and adaptive gradient methods. The convergence of SGD for smooth nonconvex functions was first analyzed in (Ghadimi and Lan, 2013), assuming a uniform bound on the variance of stochastic gradients. Arjevani et al. (2023) gave a tight lower bound. Bottou et al. (2018) extended the convergence result with the so-called affine variance noise model. Different assumptions on the second moment of stochastic gradients appearing in the literature were reviewed in (Khaled and Richtárik, 2023), with a view to developing a general convergence theory of SGD for smooth nonconvex functions.
Appendix B The Overall Algorithm
@@
$\mathsf{T}_{j}(x_{k};\mathcal{B}_{k},\eta)\coloneqq\frac{1}{b_{k}-1}\sum_{i\in\mathcal{B}_{k}}\left(\partial_{j}f(x_{k};\xi_{i})-\partial_{j}F_{\mathcal{B}_{k}}(x_{k})\right)^{2}\big{/}\left(\eta^{2}(\partial_{j}F_{\mathcal{B}_{k}}(x_{k}))^{2}\right),\;j\in\llbracket d\rrbracket$
@@
Compute the norm test statistic
$\hskip 56.9055pt\mathsf{T}\equiv\mathsf{T}(x_{k};\mathcal{B}_{k},\eta)\coloneqq\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))\big{/}\left(\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}\right)$
@@
$\hskip 42.67912pt\mathsf{T}_{\mathrm{ip}}(x_{k};\mathcal{B}_{k},\vartheta)\coloneqq\mathrm{Var}_{i\in\mathcal{B}_{k}}(\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle)\big{/}\left(\vartheta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{4}\right)$
@@
Compute the orthogonality test statistic
$\hskip 71.13188pt\mathsf{T}_{\mathrm{ortho}}(x_{k};\mathcal{B}_{k},\nu)\coloneqq\mathsf{V}(x_{k};\mathcal{B}_{k})/\left(\nu^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}\right)$
@@
[…]
Appendix C Proofs of Main Text
@@ C.1 Proof of Equation 1
Note that $-\nabla F_{\mathcal{B}}(x)$ is a descent direction for $F$ if and only if $\langle\nabla F_{\mathcal{B}}(x),\nabla F(x)\rangle\geqslant 0$. Then we have, for $\eta\in\left[0,1\right)$, for any $x\in\mathbb{R}^{d}$,
$\|\nabla F_{\mathcal{B}}(x)-\nabla F(x)\|^{2}=\|\nabla F_{\mathcal{B}}(x)\|^{2}-2\langle\nabla F_{\mathcal{B}}(x),\nabla F(x)\rangle+\|\nabla F(x)\|^{2}\leqslant\eta^{2}\|\nabla F(x)\|^{2}\\ \iff-2\langle\nabla F_{\mathcal{B}}(x),\nabla F(x)\rangle\leqslant-(1-\eta^{2})\|\nabla F(x)\|^{2}-\|\nabla F_{\mathcal{B}}(x)\|^{2}\leqslant 0,$
@@
C.2 Preparatory Definitions, Propositions and Lemmas
@@
C.2.1 Formal Statements of Proposition 5.1
@@
Proposition C.1 (Exact variance norm test).
Suppose that the condition of the exact variance norm test (4) is satisfied for every iteration $k\in\mathbb{N}^{*}$, i.e., for some constant $\eta>0$,
@@
$\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant(1+\eta^{2})\|\nabla F(x_{k})\|^{2}.$ (C.1)
@@
Proposition C.2 (Exact variance inner product test and orthogonality test).
Suppose that the conditions of the exact variance inner product test (5) and the exact variance orthogonality test (7) are satisfied for every iteration $k\in\mathbb{N}^{*}$, i.e., for some constant , i.e., for some constant $\vartheta>0$,
@@
$\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant(1+\vartheta^{2}+\nu^{2})\|\nabla F(x_{k})\|^{2}.$ (C.2)
@@
This is part of the result of (Bollapragada et al., 2018a), Lemma 3.1. We include its proof here with our notation for completeness.
@@
$\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant\frac{\mathbb{E}_{k}\left[\left(\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle\right)^{2}\right]}{\|\nabla F(x_{k})\|^{2}}+\nu^{2}\|\nabla F(x_{k})\|^{2}.$ (C.3)
It remains to bound the first term on the right hand side of (C.3), using the condition of the exact variance inner product test. Since (5) is satisfied, the following inequality holds:
@@
Along with (C.3), this yields
@@
which is identical to (4). The desired conclusion follows from Proposition C.1. ∎
C.2.3 Technical Lemmas